# Empowering Multi-Robot Cooperation via Sequential World Models

### v1

@@ C.2 Additional Behavior Visualization
Figure 19: Visualization of the learned behaviors on Gate task.
@@
Figure 20: Visualization of the learned behaviors on Shepherd task.
For the Shepherd task, from $t=3$ (Frame C) to $t=4$ (Frame D), the sheep approaches the upper boundary. At this point, Robot 1 moves downward while Robot 2 advances to prevent the sheep from escaping, effectively guiding it back to the target area. This sequence highlights SeqWM’s predictive planning and adaptive role assignment: the leading robot influences the trajectory of the target robot, while the following robot adjusts accordingly to ensure successful shepherding.
C.3 Role of Actor and Planner
@@
C.4 Early-Stopping Planner
Inference Time Cost. We report the per-step execution time of SeqWM on the BottleCap task using a single RTX A6000 GPU on the left side of Figure 21. The execution time increases almost linearly with the rollout horizon $H$ and the number of planner iterations $K$, which is consistent with the design of SeqWM. With the default settings, SeqWM achieves a per-step execution time of 12.8 ms, making it suitable for most real-time robotic tasks.
Early-Stopping Heuristic. To further enhance the efficiency of SeqWM, we introduce an early-stopping heuristic in Section 4.2 that terminates iterations when the change in the action distribution is not significant. The KL divergence is used as a measure of distribution change, and the execution time and performance under different thresholds on the BottleCap task are shown on the right side of Figure 21. When the threshold is set to 0.5, SeqWM reduces the execution time by approximately 27.3% while incurring only about 5.9% performance loss.
@@ C.5 Sim-to-Real Deployment
  • Domain randomization: Taking the PushBox task as an example, we randomize the initial positions/orientations of both robots and the box, the position and distance of the target, and the friction coefficient between the box and floor to improve robustness to environmental variations.

### v2

@@ C.3 Additional Behavior Visualization
Figure 20: Visualization of the learned behaviors on Gate.
@@
Figure 21: Visualization of the learned behaviors on Shepherd.
For the Shepherd, from $t=3$ (Frame C) to $t=4$ (Frame D), the sheep approaches the upper boundary. At this point, Robot 1 moves downward while Robot 2 advances to prevent the sheep from escaping, effectively guiding it back to the target area. This sequence highlights SeqWM’s predictive planning and adaptive role assignment: the leading robot influences the trajectory of the target robot, while the following robot adjusts accordingly to ensure successful shepherding.
C.4 Role of Actor and Planner
@@
C.5 Early-Stopping Planner
Inference Time Cost. We report the per-step execution time of SeqWM on BottleCap using a single RTX A6000 GPU on the left side of Figure 22. The execution time increases almost linearly with the rollout horizon $H$ and the number of planner iterations $K$, which is consistent with the design of SeqWM. With the default settings, SeqWM achieves a per-step execution time of 12.8 ms, making it suitable for most real-time robotic tasks.
Early-Stopping Heuristic. To further enhance the efficiency of SeqWM, we introduce an early-stopping heuristic in Section 4.2 that terminates iterations when the change in the action distribution is not significant. The KL divergence is used as a measure of distribution change, and the execution time and performance under different thresholds on BottleCap are shown on the right side of Figure 22. When the threshold is set to 0.5, SeqWM reduces the execution time by approximately 27.3% while incurring only about 5.9% performance loss.
@@ C.6 Sim-to-Real Deployment
  • Domain randomization: Taking PushBox as an example, we randomize the initial positions/orientations of both robots and the box, the position and distance of the target, and the friction coefficient between the box and floor to improve robustness to environmental variations.